# Asymptotics for conformal inference

### v1

@@ 2.1 Prediction setting
In the classical setting, the variables $Z_{i},i\in\llbracket n+m\rrbracket$, are i.i.d., hence Assumption 1 is true with $P_{\tiny\mbox{cal}}=P_{\tiny\mbox{test}}$. Since the vector of scores $\mathcal{D}_{{\tiny\mbox{cal}}}\cup\mathcal{D}_{{\tiny\mbox{test}}}$ is exchangeable in this case, the conformal $p$-values are marginally super-uniform (Vovk et al.,, 2005; Romano and Wolf,, 2005) which leads to non-asymptotically valid prediction sets. However, in case of a distribution shift between the distributions of the calibration and the test sample, that is $P_{\tiny\mbox{cal}}\neq P_{\tiny\mbox{test}}$, this property is lost. To solve this issue, Tibshirani et al., (2019); Barber et al., (2023) proposed in this case to use weighted conformal prediction. We follow this approach (with light formal variations for mathematical convenience) by introducing a nonnegative weight function $w:\mathbb{R}\rightarrow\mathbb{R}^{+}$ and $w(+\infty)\in\mathbb{R}^{+}$ and the prediction set
@@
$\mathcal{C}^{w,\alpha}(X_{n+1})\coloneqq\left\{y\in\mathcal{Y}:\ S(X_{n+1},y)\leqslant Q_{1-\alpha}\left({\frac{w(+\infty)}{W}\delta_{+\infty}+\sum_{k\in\llbracket n\rrbracket}\frac{w\left(S_{k}\right)}{W}\delta_{S_{k}}}\right)\right\},$
where $Q_{\alpha}(\mu)$ denotes the $\alpha$-quantile of the probability measure $\mu$, $\delta_{a}$ denotes the Dirac measure in $a$ and $W:=w(+\infty)+\sum_{k\in\llbracket n\rrbracket}w\left(S_{k}\right)$ is a normalization constant. The set $\mathcal{C}^{w,\alpha}(X_{n+1})$ can also be described with weighted conformal $p$-values thanks to Lemma F.1: $Y_{n+1}\notin\mathcal{C}^{w,\alpha}(X_{n+1})$ if and only if ${p_{1}^{w,(n)}}\leqslant\alpha$, with
${p_{1}^{w,(n)}}\coloneqq\frac{w(+\infty)+\sum_{k=1}^{n}w(S_{k}){\mathds{1}_{S_{k}\geqslant T_{1}}}}{w(+\infty)+\sum_{k=1}^{n}w(S_{k})}.$
Similarly, with a test sample of size $m$, we obtained the $m$ weighted conformal $p$-values,
@@
${p_{i}^{w,(n)}}\coloneqq\frac{w(+\infty)+\sum_{k=1}^{n}w(S_{k}){\mathds{1}_{S_{k}\geqslant T_{i}}}}{w(+\infty)+\sum_{k=1}^{n}w(S_{k})},\text{ for all $i\in\llbracket m\rrbracket$},$ (6)
@@
The $\mathrm{FCP}$ of these weighted conformal prediction sets is in this case
@@
which, similarly to (5), is the e.c.d.f. of the weighted conformal $p$-values family.
@@
Clearly, if the calibration and test sample are exchangeable ($P_{\tiny\mbox{cal}}=P_{\tiny\mbox{test}}$), the function $w^{*}$ is constantly equal to $1$. However, under a distribution shift ($P_{\tiny\mbox{cal}}\neq P_{\tiny\mbox{test}}$), the oracle weight function is different, and oracle-weighted conformal $p$-values ${p_{i}^{w^{*},(n)}}$ (6) are different from the regular ones ${p_{i}^{(n)}}$ (4). They recover the marginally super-uniform property provided that $w^{*}(+\infty)\geqslant\sup_{u\in\mathbb{R}}w^{*}(u)$, see Tibshirani et al., (2019).
@@ 2.2 Novelty detection setting
In the novelty detection setting (see for instance Vovk et al.,, 2005; Bates et al.,, 2023), we observe a calibration set $\{Z_{k},k\in\llbracket n\rrbracket\}$ of size $n$, with values in $\mathcal{Z}$ distributed according to an unknown “null” distribution $\mathbb{P}_{0}$ and a test sample $\{Z_{n+i},i\in\llbracket m\rrbracket\}$ with $Z_{n+i}\in\mathcal{Z}$ either distributed as $\mathbb{P}_{0}$ or not. Formally, we introduce a subset ${\mathcal{H}}_{0}\subset\{1,2,\dots\}$ so that $Z_{n+i}\sim\mathbb{P}_{0}$ when $i\in{\mathcal{H}}_{0}$. We also denote ${\mathcal{H}}_{1}=\{1,2,\dots\}\backslash{\mathcal{H}}_{0}$. In addition, it is given a non-conformity score function $S:\mathcal{Z}\rightarrow\mathbb{R}$ such that $S(Z)\in\mathbb{R}$ measures the non-conformity of the variables $Z$ with respect to $\mathbb{P}_{0}$ and we denote $\mathcal{D}_{{\tiny\mbox{cal}}}=\left\{S_{k}:k\in\llbracket n\rrbracket\right\}$ and $\mathcal{D}_{{\tiny\mbox{test}}}=\left\{T_{i}:i\in\llbracket m\rrbracket\right\}$ the set of the score from the calibration and test samples respectively. A novelty detection procedure decides, for each $i\in\llbracket m\rrbracket$, whether $Z_{n+i}$ is a novelty (that is, does not follow $\mathbb{P}_{0}$), or not. The procedure described by Bates et al., (2023) consists in computing the conformal $p$-values defined in (4) (by using the specific $S_{k}$’s and $T_{i}$’s of novelty detection), and then applying the Benjamini-Hochberg procedure (Benjamini and Hochberg,, 1995) on this $p$-value family $({p_{i}^{(n)}})_{i\in\llbracket m\rrbracket}$ (see Section 5.2.1 for more formal details). This gives a rejection set $\mathcal{R}_{\alpha}\subset\llbracket m\rrbracket$ corresponding to the indices of the declared novelties.
$\mathrm{FDP}_{m}^{(n)}(\mathcal{R}_{\alpha})$ $\coloneqq\frac{\left|\mathcal{R}_{\alpha}\cap{\mathcal{H}}_{0}\right|}{\left|\mathcal{R}_{\alpha}\right|\vee 1};$ (9)
$\mathrm{TDP}_{m}^{(n)}(\mathcal{R}_{\alpha})$ $\coloneqq\frac{\left|\mathcal{R}_{\alpha}\cap{\mathcal{H}}_{1}\right|}{\left|{\mathcal{H}}_{1}\right|\vee 1}.$ (10)
The $\mathrm{FDP}$ corresponds to the proportion of errors among the declared novelties (related to a type I error notion), while the $\mathrm{TDP}$ corresponds to the proportion of correct decisions among the true novelties (related to a power notion). The following will be assumed throughout the paper for (weighted or not) novelty detection:
@@
The set of calibration scores $\left(S_{k}\right)_{k\geqslant 1}$ and the set of test scores $\left(T_{i}\right)_{i\geqslant 1}$ are two independent families of real random variables. The variables $S_{k}$, $k\geqslant 1$, are i.i.d. with distribution $P_{\tiny\mbox{cal}}$ and c.d.f. $F_{\tiny\mbox{cal}}$. The variables $T_{i}$, $i\geqslant 1$, are independent, the variables $T_{i}$, $i\in{\mathcal{H}}_{0}$, are identically distributed as a null score distribution $P_{0}$ and c.d.f. $F_{0}$, and the variables $T_{i}$, $i\in{\mathcal{H}}_{1}$, are identically distributed as an alternative score distribution $P_{\tiny\mbox{test}}$ (potentially different from $P_{0}$) with c.d.f. $F_{\tiny\mbox{test}}$. Moreover, $F_{\tiny\mbox{cal}}$, $F_{0}$ and $F_{\tiny\mbox{test}}$ are continuous.
We also consider the case of a distribution shift between the distribution of the calibration set $P_{\tiny\mbox{cal}}$ and the null distribution $P_{0}$, in which case we propose to use a weighted $p$-value approach (as for the prediction task). For some weight function $w:\mathbb{R}\rightarrow\mathbb{R}^{+}$ and $w(+\infty)\in\mathbb{R}^{+}$, the weighted conformal $p$-values are the ones from (6) (by using the specific $S_{k}$’s and $T_{i}$’s of novelty detection). We introduce the weighted Bates et al., (2023) procedure as the one applying the BH procedure on this weighted $p$-value family and we denote its rejection set by $\mathcal{R}^{w}_{\alpha}\subset\llbracket m\rrbracket$. We defined as above the $\mathrm{FDP}$ and $\mathrm{TDP}$ of this procedure.
@@
The oracle weighted conformal $p$-values ${p_{i}^{w^{*},(n)}}$, $i\in{\mathcal{H}}_{0}$, are marginally super-uniform provided that $w^{*}(+\infty)\geqslant\sup_{u\in\mathbb{R}}w^{*}(u)$, see Tibshirani et al., (2019).
@@ 2.3 Spaces for process convergence
We study the asymptotic convergence of random processes and we consider usual spaces defined in Billingsley, (1999) and van der Vaart and Wellner, (1996), denoted as usual $D[0,1]$, $D(0,1)$ and $\ell^{\infty}(0,1)$, and which are briefly described below.
First, $D[0,1]$ is the set of càdlàg function $f:[0,1]\rightarrow\mathbb{R}$ with the usual Skorohod topology. Second, $D(0,1)$ is the set of càdlàg function $f:(0,1)\rightarrow\mathbb{R}$ with the extended Skorohod topology [^2] This definition is similar to the convergence in $D[0,+\infty)$ considered in Billingsley, (1999) defined as follows: $(x_{n})_{n}\in D(0,1)^{\mathbb{N}}$ converge to $x\in D(0,1)$ if and only if, for all $(a,b)\in(0,1)^{2}$ with $a<b$ such that $x$ is continuous at points $a$ and $b$, the sequence $(x_{n})_{n}$ restricted to $[a,b]$ converges to $x$ restricted to $[a,b]$ in $D[a,b]$ with the usual Skorohod topology. Finally, $\ell^{\infty}(0,1)$ is the space of locally bounded function $f:(0,1)\rightarrow\mathbb{R}$ with the topology of the uniform convergence on all compact sets of $(0,1)$: a sequence $(f_{n})_{n}\in[\ell^{\infty}(0,1)]^{\mathbb{N}}$ converges to $f\in\ell^{\infty}(0,1)$ if and only if for all $K$ compact subsets of $(0,1)$, $\sup_{x\in K}\left|f_{n}(x)-f(x)\right|\rightarrow 0$. We also consider $D(\mathbb{R})$ the space of càdlàg function from $\mathbb{R}$ to $\mathbb{R}$ with the extended Skorohod topology and $\ell^{\infty}(\mathbb{R})$ the space of locally bounded function from $\mathbb{R}$ to $\mathbb{R}$ with the topology of the uniform convergence on all compact sets of $\mathbb{R}$.

### v3

@@ 2.1 Prediction setting
When the variables $Z_{i},i\in\llbracket n+m\rrbracket$, are i.i.d., Assumption 1 is true with $P_{\tiny\mbox{cal}}=P_{\tiny\mbox{test}}$ conditionally on $\mathcal{D}_{{\tiny\mbox{train}}}$, see Remark 2.1 below for more details. Since the vector of scores $\mathcal{D}_{{\tiny\mbox{cal}}}\cup\mathcal{D}_{{\tiny\mbox{test}}}$ is exchangeable in this case, the conformal $p$-values are marginally super-uniform (Vovk, Gammerman and Shafer, 2005; Romano and Wolf, 2005) which leads to non-asymptotically valid prediction sets. However, in case of a distribution shift between the distributions of the calibration and the test sample, that is $P_{\tiny\mbox{cal}}\neq P_{\tiny\mbox{test}}$, this property is lost. To solve this issue, Tibshirani et al. (2019); Barber et al. (2023) proposed in this case to use weighted conformal prediction. We follow this approach (with light formal variations for mathematical convenience) by introducing a nonnegative weight function $w:\mathbb{R}\rightarrow\mathbb{R}^{+}$, $w(+\infty)\in\mathbb{R}^{+}$ and the prediction set
@@
$\mathcal{C}^{w,\alpha}(X_{n+1})\coloneqq\mathopen{}\mathclose{{\left\{y\in\mathcal{Y}:\ S(X_{n+1},y)\leqslant Q_{1-\alpha}\mathopen{}\mathclose{{\left({\frac{w(+\infty)}{W}\delta_{+\infty}+\sum_{k\in\llbracket n\rrbracket}\frac{w\mathopen{}\mathclose{{\left(S_{k}}}\right)}{W}\delta_{S_{k}}}}}\right)}}\right\},$
where $Q_{1-\alpha}(\mu)$ denotes the $1-\alpha$-quantile of the probability measure $\mu$, $\delta_{a}$ denotes the Dirac measure in $a$ and $W:=w(+\infty)+\sum_{k\in\llbracket n\rrbracket}w\mathopen{}\mathclose{{\left(S_{k}}}\right)$ is a normalization constant. There are different ways to choose the weight function $w$. In the case of a known distribution shift in some specific model, the weight function is usually chosen as the Radon-Nikodym derivative of the shift (Tibshirani et al., 2019; Jin and Candès, 2023). Barber et al. (2023) assumes that the weight function $w$ is deterministic and fixed by the user before looking at the data. To accommodate both points of view, we assume that $w$ can be estimated from a (possibly infinite) independent training dataset (potentially different from $\mathcal{D}_{{\tiny\mbox{train}}}$ above) containing both data with the same distribution as the calibration points, and data with the same distribution as the test points. Our results hold conditionally on this training sample and $w$ can be considered as deterministic. Thanks to Lemma F.1, the set $\mathcal{C}^{w,\alpha}(X_{n+1})$ can be described with weighted conformal $p$-values: $Y_{n+1}\notin\mathcal{C}^{w,\alpha}(X_{n+1})$ if and only if ${p_{1}^{w,(n)}}\leqslant\alpha$, with $p_{1}^{w,(n)}$} the (unobserved) weighted conformal $p$-value given by (7) below. Similarly, with a test sample of size $m$, we obtained $m$ weighted conformal $p$-values,
@@
${p_{i}^{w,(n)}}\coloneqq\frac{w(+\infty)+\sum_{k=1}^{n}w(S_{k}){\mathds{1}_{S_{k}\geqslant T_{i}}}}{w(+\infty)+\sum_{k=1}^{n}w(S_{k})},\text{ for all $i\in\llbracket m\rrbracket$}.$ (7)
@@
The $\mathrm{FCP}$ of these weighted conformal prediction sets is then defined by
@@
which, similarly to (6), is the e.c.d.f. of the weighted conformal $p$-values family. As for the unweighted case, the weighted conformal $p$-values are not observed, and neither is the corresponding false coverage proportion.
@@
Clearly, if the calibration and test sample are exchangeable ($P_{\tiny\mbox{cal}}=P_{\tiny\mbox{test}}$), the function $w^{*}$ is constantly equal to $1$. However, under a distribution shift ($P_{\tiny\mbox{cal}}\neq P_{\tiny\mbox{test}}$), the oracle weight function is different, and oracle-weighted conformal $p$-values $p_{i}^{w^{*},(n)}$} (7) are different from the regular ones $p_{i}^{(n)}$} (4). They satisfy the marginally super-uniform property provided that $w^{*}(+\infty)\geqslant\sup_{u\in\mathbb{R}}w^{*}(u)$, see Tibshirani et al. (2019).
Remark 2.1.
In split conformal inference the score function $S(\cdot)$ and the weight function $w(\cdot)$ are constructed here with an independent training dataset which creates a dependence among the scores $(S(Z_{i}))_{i\in\llbracket n+m\rrbracket}$. However, our results are stated conditionally on the training phase hence the weight function $w(\cdot)$ and score function $S(\cdot)$ can be considered as deterministic and the scores $(S(Z_{i}))_{i\in\llbracket n+m\rrbracket}$ are independent as stated in Assumption 1 provided that the data $(Z_{i})_{i\in\llbracket n+m\rrbracket}$ are independent (which is a common assumption in the conformal prediction literature).
@@ 2.2 Novelty detection setting
In the novelty detection setting (see for instance Vovk, Gammerman and Shafer, 2005; Bates et al., 2023), we observe a calibration set $\{Z_{k},k\in\llbracket n\rrbracket\}$ of size $n$, with values in $\mathcal{Z}$ distributed according to an unknown “null” distribution $\mathbb{P}_{0}$ and a test sample $\{Z_{n+i},i\in\llbracket m\rrbracket\}$ with $Z_{n+i}\in\mathcal{Z}$ either distributed as $\mathbb{P}_{0}$ or not. Formally, we introduce a subset ${\mathcal{H}}_{0}\subset\{1,2,\dots\}$ so that $Z_{n+i}\sim\mathbb{P}_{0}$ when $i\in{\mathcal{H}}_{0}$. We also denote ${\mathcal{H}}_{1}=\{1,2,\dots\}\backslash{\mathcal{H}}_{0}$. We assume that we have access to an independent training dataset $\mathcal{D}_{{\tiny\mbox{train}}}$ used to learn a non-conformity score function $S:\mathcal{Z}\rightarrow\mathbb{R}$ such that $S(Z)\in\mathbb{R}$ measures the non-conformity of the variables $Z$ with respect to $\mathbb{P}_{0}$. Our results are stated conditionally on $\mathcal{D}_{{\tiny\mbox{train}}}$ and the training phase. Therefore, as in the prediction setting, we assume that the score function $S(\cdot)$ is fixed. We denote $\mathcal{D}_{{\tiny\mbox{cal}}}=\mathopen{}\mathclose{{\left\{S_{k}\,:\,k\in\llbracket n\rrbracket}}\right\}$ for $S_{k}\coloneqq S(Z_{k})$, and $\mathcal{D}_{{\tiny\mbox{test}}}=\mathopen{}\mathclose{{\left\{T_{i}\,:\,i\in\llbracket m\rrbracket}}\right\}$ for $T_{i}:=S(Z_{n+i})$ the set of the scores from the calibration and test samples respectively. A novelty detection procedure decides, for each $i\in\llbracket m\rrbracket$, whether $Z_{n+i}$ is a novelty (that is, does not follow $\mathbb{P}_{0}$), or not. The procedure described by Bates et al. (2023) consists in computing the conformal $p$-values defined in (4) (by using the specific $S_{k}$’s and $T_{i}$’s of novelty detection which are observed in this case), and then applying the Benjamini-Hochberg procedure (Benjamini and Hochberg, 1995) on this $p$-value family $({p_{i}^{(n)}})_{i\in\llbracket m\rrbracket}$ (see Section 5.2.1 for more formal details). This gives a rejection set $\mathcal{R}_{\alpha}\subset\llbracket m\rrbracket$ corresponding to the indices of the declared novelties. We introduce the two main quantities to measure the performance of the rejection set,
$\mathrm{FDP}_{m}^{(n)}(\mathcal{R}_{\alpha})$ $\coloneqq\frac{\mathopen{}\mathclose{{\left|\mathcal{R}_{\alpha}\cap{\mathcal{H}}_{0}}}\right\rvert}{\mathopen{}\mathclose{{\left|\mathcal{R}_{\alpha}}}\right\rvert\vee 1};$ (10)
$\mathrm{TDP}_{m}^{(n)}(\mathcal{R}_{\alpha})$ $\coloneqq\frac{\mathopen{}\mathclose{{\left|\mathcal{R}_{\alpha}\cap{\mathcal{H}}_{1}}}\right\rvert}{\mathopen{}\mathclose{{\left|{\mathcal{H}}_{1}}}\right\rvert\vee 1}.$ (11)
The $\mathrm{FDP}$ corresponds to the proportion of errors among the declared novelties (related to a type I error notion), while the $\mathrm{TDP}$ corresponds to the proportion of correct decisions among the true novelties (related to power and type II error notions). In the novelty detection setting all the scores are available hence the conformal $p$-values can be computed. Nevertheless, the $\mathrm{FDP}$ and the $\mathrm{TDP}$ remain unknown since the indices of the nulls/alternatives are not observed. The following is assumed throughout the paper for (weighted) novelty detection.
@@
The set of calibration scores $\mathopen{}\mathclose{{\left(S_{k}}}\right)_{k\geqslant 1}$ and the set of test scores $\mathopen{}\mathclose{{\left(T_{i}}}\right)_{i\geqslant 1}$ are two independent families of real random variables. The variables $S_{k}$, $k\geqslant 1$, are i.i.d. with distribution $P_{\tiny\mbox{cal}}$ and c.d.f. $F_{\tiny\mbox{cal}}$. The variables $T_{i}$, $i\geqslant 1$, are independent, the variables $T_{i}$, $i\in{\mathcal{H}}_{0}$, are identically distributed as a null score distribution $P_{0}$ and c.d.f. $F_{0}$, and the variables $T_{i}$, $i\in{\mathcal{H}}_{1}$, are identically distributed as an alternative score distribution $P_{\tiny\mbox{test}}$ (potentially different from $P_{0}$) with c.d.f. $F_{\tiny\mbox{test}}$. Moreover, $F_{\tiny\mbox{cal}}$, $F_{0}$ and $F_{\tiny\mbox{test}}$ are continuous.
We also consider the case of a distribution shift between the distribution of the calibration set $P_{\tiny\mbox{cal}}$ and the null distribution $P_{0}$, in which case we propose to use a weighted $p$-value approach (as for the prediction task). The weights are assumed to be built with an independent training sample so they are considered as fixed here (our results hold conditionally on this sample) considering oracle weights as a particular case with perfect training. For some weight function $w:\mathbb{R}\rightarrow\mathbb{R}^{+}$ and $w(+\infty)\in\mathbb{R}^{+}$, the weighted conformal $p$-values are the ones from (7) (by using the specific $S_{k}$’s and $T_{i}$’s of novelty detection). We define the weighted version of the procedure of Bates et al. (2023) as the BH procedure applied to this weighted $p$-value family and we denote its rejection set by $\mathcal{R}^{w}_{\alpha}\subset\llbracket m\rrbracket$. The $\mathrm{FDP}$ and $\mathrm{TDP}$ of $\mathcal{R}^{w}_{\alpha}$ are defined similarly to above.
@@
The oracle weighted conformal $p$-values $p_{i}^{w^{*},(n)}$}, $i\in{\mathcal{H}}_{0}$, are marginally super-uniform provided that $w^{*}(+\infty)\geqslant\sup_{u\in\mathbb{R}}w^{*}(u)$, see Tibshirani et al. (2019).
Remark 2.2.
The score function $S(\cdot)$ and the weight function $w(\cdot)$ are trained here with an independent training dataset which induces a non-trivial dependence between the scores $(S(Z_{i}))_{i\in\llbracket n+m\rrbracket}$. However, our results are stated conditionally on the training phase and the score function $S(\cdot)$ and weight function $w(\cdot)$ can be considered as deterministic. Hence, the scores $(S(Z_{i}))_{i\in\llbracket n+m\rrbracket}$ satisfy Assumption 2 conditionally on the training phase as long as the underlying data $(Z_{i})_{i\in\llbracket n+m\rrbracket}$ are independent and have the same distribution within a group (calibration, null, alternative), which is a classical assumption (Bates et al., 2023; Jin and Candès, 2023).
@@ 2.3 Spaces for process convergence
We study the asymptotic convergence of random processes and we consider usual spaces defined in Billingsley (1999) and van der Vaart and Wellner (1996), denoted as usual $D[0,1]$, $D(0,1)$ and $\ell^{\infty}(0,1)$, and which are briefly described below.
First, $D[0,1]$ is the set of càdlàg functions $f:[0,1]\rightarrow\mathbb{R}$ with the usual Skorohod topology. Second, $D(0,1)$ is the set of càdlàg functions $f:(0,1)\rightarrow\mathbb{R}$ with the extended Skorohod topology [^2] This definition is similar to the convergence in $D[0,+\infty)$ considered in Billingsley (1999) defined as follows: $(x_{n})_{n}\in D(0,1)^{\mathbb{N}}$ converges to $x\in D(0,1)$ if and only if, for all $(a,b)\in(0,1)^{2}$ with $a<b$ such that $x$ is continuous at points $a$ and $b$, the sequence $(x_{n})_{n}$ restricted to $[a,b]$ converges to $x$ restricted to $[a,b]$ in $D[a,b]$ with the usual Skorohod topology. Finally, $\ell^{\infty}(0,1)$ is the space of locally bounded functions $f:(0,1)\rightarrow\mathbb{R}$ with the topology of the uniform convergence on all compact sets of $(0,1)$: a sequence $(f_{n})_{n}\in[\ell^{\infty}(0,1)]^{\mathbb{N}}$ converges to $f\in\ell^{\infty}(0,1)$ if and only if for all $K$ compact subsets of $(0,1)$, $\sup_{x\in K}\mathopen{}\mathclose{{\left|f_{n}(x)-f(x)}}\right\rvert\rightarrow 0$. We also consider $D(\mathbb{R})$ the space of càdlàg functions from $\mathbb{R}$ to $\mathbb{R}$ with the extended Skorohod topology and $\ell^{\infty}(\mathbb{R})$ the space of locally bounded functions from $\mathbb{R}$ to $\mathbb{R}$ with the topology of the uniform convergence on all compact sets of $\mathbb{R}$.